### 19 Jan 2024 v1

@@ Abstract
Quantum computing (QC) and machine learning (ML), taken individually or combined into quantum-assisted ML (QML), are ascending computing paradigms whose calculations come with huge potential for speedup, increase in precision, and resource reductions. Likely improvements for numerical simulations in engineering imply the possibility of a strong economic impact on the manufacturing industry. In this project report, we propose a framework for a quantum computing-enhanced service ecosystem for simulation in manufacturing, consisting of various layers ranging from hardware to algorithms to service and organizational layers. In addition, we give insight into the current state of the art of applications research based on QC and QML, both from a scientific and an industrial point of view. We further analyse two high-value use cases with the aim of a quantitative evaluation of these new computing paradigms for industrially-relevant settings.
@@ 1 Introduction
Two use cases are used to illustrate the opportunities and challenges of the practical application of QC. The first use case relates to the simulation of manufacturing processes with the aid of FEM for the example of the production of blade integrated disks (blisks) for jet engines. This involves the simulation and optimization of milling processes while complying with quality specifications, i.e., it has to be ensured that the milling process is not compromised by resonant mode excitations. In this use case, QC technologies are implemented to accelerate classical simulation processes. The second use case relates to the field of laser cutting of metals with a focus on heat dispersion and thermal expansion. Here, hybrid and pure QML approaches are discussed. Both cases are attractive for QC – both in the scientific computing and the QML context. They deal with large sets of data but a relatively small set of input and output information. Also, they are based on linear or near-linear theories.
@@ 3.1 State of the art
Most quantum algorithms for linear algebra rely on the quantum phase estimation (QPE) subroutine kitaevbook ; nielsenChuang ; 1998_cleve . The problem that QPE solves can be formulated completely classically: given the eigenvector $\vec{u}$ with eigenvalue $e^{i\varphi}$ of a $N\times N$ unitary matrix $U$, estimate the phase $\varphi$ of the eigenvalue. It was shown in Ref. 1999_abrams how QPE can be turned into an eigenvalue and eigenvector finder for a general matrix $H$ with an exponential reduction in number of systems needed, given some assumptions on the matrix $H$. We note that QPE is the basic subroutine used by Shor’s factoring algorithm shor1997 and by the standard version of the Harrow-Hassidim-Lloyd (HHL) algorithm for solving linear systems of equations 2009_harrow .
@@
The core idea of quantum algorithms for linear algebra is to encode a $N\times N$ matrix as an operator in a quantum system involving only $\mathcal{O}(\mathrm{poly}(\log N))$ qubits. Despite this exponential compression, a quantum computer is still capable of extracting useful information about the matrix, and thus, the problem we want to solve. However, an end-to-end analysis of a quantum algorithm requires studying the scaling also with respect to other parameters such as the sparsity or the rank of the matrix, its condition number, and the desired precision of the solution. Depending on the problem, these parameters can also implicitly depend on the size of the matrix. Additionally, the encoding step in the quantum computer and the extraction of the information from the quantum computer, which are problem-dependent tasks, also need to be efficient. We refer the reader to Ref. dalzell2023 for a general survey of relevant quantum algorithms and their scaling.
@@
An important subroutine that is employed in many quantum algorithms is Hamiltonian simulation. In Hamiltonian simulation, we are given a Hermitian matrix $H$, which we call a Hamiltonian, and a quantum system in a state |$\ket{\psi}$. The task is to transform the state into a new state evolved for a time $t$, |$\ket{\psi(t)}=e^{-iHt}\ket{\psi}$, within a certain accuracy. The first quantum algorithm for Hamiltonian simulation was given by Lloyd for the case in which the Hamiltonian has an efficient decomposition in terms of operators acting locally on qubits lloyd1996 . Later, Aharonov and Ta-Shma extended this result to sparse Hamiltonians with oracle access to their matrix elements aharonov2003 . For this access model, improvements in terms of runtime and number of ancilla systems needed have been achieved during the years 2007_Berry ; berry2012 ; Childs2010 . In recent years, the seminal works of Low and Chuang lowChuang2017 ; lowChuang2019 allowed to formulate the previous algorithms under the general framework of block-encoding of matrices into unitaries. The block-encoded formulation of Hamiltonian simulation further sparked the development of the quantum singular value transform (QSVT) gilyen2019 ; tang2023 , which, in the modern theory of quantum algorithms, is viewed as a fundamental primer of several quantum algorithms with a provable quantum advantage martyn2021 . On the other hand, the QSVT formulation of quantum algorithms has also allowed the inverse process of dequantization, i.e., the development of classical algorithms inspired by the quantum ones tang2018 ; chia2020 ; gilyenLloyd2018 ; tang2021 ; bakshi2023 ; gharibian2022 .
@@
In our project, we explore the application of quantum algorithms to problems that emerge in numerical simulations of manufacturing processes. A typical example of such a problem is the analysis of the vibrations of a solid when subject to an external force. The problem translates into an eigenvalue problem that is discretized using finite element methods (FEM) for instance. Notice that once discretized the problems become mathematically equivalent to that of coupled harmonic oscillators, which has been recently studied in Ref. babbush2023 . Quantum algorithms for FEM simulations have already been studied in Refs. clader2013 ; montanaroPallister2016 , where the authors analyzed algorithms based on the HHL algorithm 2009_harrow . General quantum algorithms for partial differential equations have also been recently investigated Berry2017 ; Childs2020 ; Childs2021highprecision ; liu2021 . We remark that QPE-based algorithms are believed to require a fault-tolerant quantum computer given their high circuit depth, and as such, they are not considered NISQ algorithms.
@@ 3.2 General approach in QUASIM
State-of-the-art manufacturing processes can be optimized with the support of computer simulations, which are time-consuming and expensive. In QUASIM, we are concerned with mathematical problems that occur frequently in industrial simulations, and quantum algorithms that solve those exact problems either in a shorter time or using less storage than any classical algorithm. We are not concentrating on improvements in absolute numbers but on its scaling with problem dependent parameters such as the number of equations to solve or the required accuracy. Quantum algorithms such as the HHL algorithm or Shor’s algorithm promise exponential speedup shor1997 ; 2009_harrow and with it favor large problems. Hence, we consider computations of large mathematical problems commonly used in simulations.
One common simulation algorithm used in manufacturing is the FEM. This method allows to solve physical equations for continuous systems in return for extensive computational effort. As part of FEM, the physical domain is discretized in so-called elements. The interaction between those elements is described by a system of numerical equations, which often require iterative solvers. For linear equations, both, the computation time ($\mathcal{O}(N^{3})$) and the required storage ($\mathcal{O}(N^{2})$ ) scales polynomially with the number of equations $N$ krishnamoorthy_matrix_2013 . The interested reader will find a more detailed description of the finite element method in Refs. reddy2005 ; huebner1975 .
In order to achieve an advantage, we concentrate on problems that can be solved with exponential speed-up over their classical counterparts. Two quantum algorithms promising such a speed-up are the HHL algorithm and the QPE, which solve linear equations and eigenvalue problems, respectively. Both problem types can be found in finite element simulations used in manufacturing. Linear equation solvers are more versatile, which makes an end-to-end implementation of the HHL algorithm more favorable. However, since QPE is a significant subroutine of the HHL algorithm, every knowledge gained in an end-to-end implementation of the QPE can be used for the HHL algorithm as well. This is why we concentrate our work on the QPE and eigenvalue problems in QUASIM.
When using QPE, one has to be cautious not to lose any speed-up in the subroutines such as the Hamiltonian simulation, state preparation, or post-processing steps. Those subroutines are the interface between classical and quantum computations and scale with the amount of information that needs to enter and exit the quantum computer. Hence, the amount of data that shall be transferred between those two platforms has to be minimal in order to contain quantum speed-up ($\lesssim\mathcal{O}(N)$ for polynomial or $\lesssim\mathcal{O}(\log N)$ for exponential speed up). Furthermore, the QPE is a probabilistic algorithm, which means that some results can only be extracted by sampling multiple times from the algorithm.
@@ 3.3 Specific application: milling dynamics simulation
Within the digital twin framework of the Fraunhofer Institute for Productions Technology dPart® ganser2021 , the engagement simulation is a dexel-based simulation. A dexel model is a discrete representation of a solid volume based on a regular grid, which enables a more efficient simulation of material removal processes than mesh-based methods. For a good representation of 3D volumes, a multi-dexel approximation is used stautner2006simulation . The engagement simulation provides the tool’s contact area, contact length, and width (see Fig. 2). It also allows tracking removed volume to enable an evaluation of the workpiece cut by cut. To compute the removed material, the algorithm calculates the intersections between the milling tool volume and the three-dimensional dexel grid. The intersections are performed individually for each of the three dexel grids to take advantage of memory closeness. The calculations performed by the intersectors are based on floating-point calculations.
@@
Starting after the dexel-based engagement simulation in CAM, the main steps of the FEM simulation workflow are: IPW conversion from dexel to solid, meshing process, and modal analysis. Considering the condition for quantum acceleration, the workflow is analyzed step by step both in terms of computational time and with regard to extracting numerical problems that can be transformed in quantum algorithms. The “modal analysis”, i.e., solving the eigenvalue problem, is found to be a suitable numerical problem for computation on a quantum computer. The solution of the eigenvalue problem using quantum algorithms is under investigation within this project. Within the classical step “modal analysis” the equation of motion
@@ 4.1 State of the art
Since NISQ computers are still in their infancy, and limited in size and available gate depths, they often can be simulated on classical hardware. Therefore one can use real and simulated NISQ computers alongside classical computers to perform tests on machine learning tasks in hybrid QC. A popular implementation of this uses variational or parameterised quantum circuits (PQCs), which have an advantage over classical ML in certain learning tasks. bowles2023contextuality .
@@ 4.2 General approach in QUASIM
Figure 3: The PQC $M_{d}$ used to make temperature predictions in the laser cutting use case. The first qubit, represents $v$ and the subsequent qubits represent the $d$ neigboring nodes. The temperature of each node is embedded on to the corresponding qubit. This is followed by a parameterised layer, where the first qubit and the subsequent qubits are entangled. Finally, the expectation value of the first qubit is measured and then projected to a temperature value.
@@ 4.3 Specific application: laser cutting
Accurately simulating the temperature during laser cutting with complex geometries and large cutting times requires significant computational efforts. In contrast to conventional FEM-based simulations, ML approaches, such as MeshGraphNets pfaff2020mgn , accelerate FEM simulations by one to two orders of magnitude whilst maintaining a high accuracy. However, training of such large ML models currently takes several weeks, which makes updating such models computationally challenging. In this work, the advantages of quantum machine learning approaches, as described above, are investigated to simulate thermodynamics from laser-induced heat on metal plates.
The metal geometry is represented through a mesh structure which can be formulated as a graph $G=(V,E)$. The model is formulated employing the PQC approach outlined in section 4.2 for predicting the temperature $f_{v}^{t+1}$ at the vertex $v$ for the next time step $(t+1)$, depending on the temperatures from the previous time step $t$ at the vertex $v$ and its $d$ neighbors $\mathcal{N}(v)$. A visualisation of the circuit can be seen in Fig. 3.

### 31 Jan 2024 v2

@@ Abstract
Quantum computing (QC) and machine learning (ML), taken individually or combined into quantum-assisted ML (QML), are ascending computing paradigms whose calculations come with huge potential for speedup, increase in precision, and resource reductions. Likely improvements for numerical simulations in engineering imply the possibility of a strong economic impact on the manufacturing industry. In this project report, we propose a framework for a quantum computing-enhanced service ecosystem for simulation in manufacturing, consisting of various layers ranging from hardware to algorithms to service and organizational layers. In addition, we give insight into the current state of the art of applications research based on QC and QML, both from a scientific and an industrial point of view. We further analyze two high-value use cases with the aim of a quantitative evaluation of these new computing paradigms for industrially-relevant settings.
@@ 1 Introduction
Two use cases are used to illustrate the opportunities and challenges of the practical application of QC. The first use case relates to the simulation of manufacturing processes with the aid of FEM for the example of the production of blade integrated disks (blisks) for jet engines. This involves the simulation and optimization of milling processes while complying with quality specifications, i.e., it has to be ensured that the milling process is not compromised by resonant mode excitations. In this use case, we investigate how quantum agorithms accelerate classical simulation processes. The second use case relates to the field of laser cutting of metals with a focus on heat dispersion and thermal expansion. Here, hybrid and pure QML approaches are discussed. Both cases are attractive for QC – both in the scientific computing and the QML context. They deal with large sets of data but a relatively small set of input and output information. Also, they are based on linear or near-linear theories.
@@ 3.1 State of the art
Most quantum algorithms for linear algebra rely on the quantum phase estimation (QPE) subroutine kitaevbook ; nielsenChuang ; 1998_cleve . The problem that QPE solves can be formulated completely classically: given the eigenvector $\vec{u}$ with eigenvalue $e^{i\varphi}$ of a $N\times N$ unitary matrix $U$, estimate the phase $\varphi$ of the eigenvalue. It was shown in Ref. 1999_abrams how QPE can be turned into an eigenvalue and eigenvector finder for a general matrix $H$ with an exponential reduction in the number of systems needed, given some assumptions on the matrix $H$. We note that QPE is the basic subroutine used by Shor’s factoring algorithm shor1997 and by the standard version of the Harrow-Hassidim-Lloyd (HHL) algorithm for solving linear systems of equations 2009_harrow .
@@
The core idea of quantum algorithms for linear algebra is to encode a $N\times N$ matrix as an operator in a quantum system involving only $\mathcal{O}(\mathrm{poly}(\log N))$ qubits. Despite this exponential compression, a quantum computer is still capable of extracting useful information about the matrix, and thus, the problem we want to solve. However, an end-to-end analysis of a quantum algorithm requires studying the scaling also with respect to other parameters such as the sparsity or the rank of the matrix, its condition number, and the desired precision of the solution. Depending on the problem, these parameters can also implicitly depend on the size of the matrix. Additionally, the encoding step in the quantum computer and the extraction of the information from the quantum computer, which are problem-dependent tasks, also need to be efficient. For a comprehensive overview of quantum algorithms and their scaling, please refer to Ref. dalzell2023 .
@@
An important subroutine that is employed in many quantum algorithms is Hamiltonian simulation. In Hamiltonian simulation, we are given a Hermitian matrix $H$, which we call a Hamiltonian, and a quantum system in a state |$\ket{\psi}$. The task is to transform the state into a new state evolved for a time $t$, |$\ket{\psi(t)}=e^{-iHt}\ket{\psi}$, within a certain accuracy. The first quantum algorithm for Hamiltonian simulation was given by Lloyd for the case in which the Hamiltonian has an efficient decomposition in terms of operators acting locally on qubits lloyd1996 . Later, Aharonov and Ta-Shma extended this result to sparse Hamiltonians with oracle access to their matrix elements aharonov2003 . For this access model, improvements in terms of runtime and number of ancilla systems needed have been achieved during the years 2007_Berry ; berry2012 ; Childs2010 . In recent years, the seminal works of Low and Chuang lowChuang2017 ; lowChuang2019 allowed the formulation of the previous algorithms under the general framework of block-encoding of matrices into unitaries. The block-encoded formulation of Hamiltonian simulation further sparked the development of the quantum singular value transform (QSVT) gilyen2019 ; tang2023 , which, in the modern theory of quantum algorithms, is viewed as a fundamental primer of several quantum algorithms with a provable quantum advantage martyn2021 . Ref. Kikuchi2023 even provided a small scale test of Hamiltonian simulation based on block encoding on NISQ devices. On the other hand, the QSVT formulation of quantum algorithms has also allowed the inverse process of dequantization, i.e., the development of classical algorithms inspired by the quantum ones tang2018 ; chia2020 ; gilyenLloyd2018 ; tang2021 ; bakshi2023 ; gharibian2022 .
@@
In our project, we explore the application of quantum algorithms to problems that emerge in numerical simulations of manufacturing processes. A typical example of such a problem is the analysis of the vibrations of a solid when subject to an external force. The problem translates into an eigenvalue problem that is discretized using finite element methods (FEM), for instance. Notice that once discretized the problem becomes mathematically equivalent to that of coupled harmonic oscillators, which has been recently studied in Ref. babbush2023 . Quantum algorithms for FEM simulations have already been studied in Refs. clader2013 ; montanaroPallister2016 , where the authors analyzed algorithms based on the HHL algorithm 2009_harrow . General quantum algorithms for partial differential equations have also been recently investigated Berry2017 ; Childs2020 ; Childs2021highprecision ; liu2021 . We remark that QPE-based algorithms are believed to require a fault-tolerant quantum computer given their high circuit depth, and as such, they are not considered NISQ algorithms.
@@ 3.2 General approach in QUASIM
State-of-the-art manufacturing processes can be optimized with the support of computer simulations, which are time-consuming and expensive. In QUASIM, we are concerned with mathematical problems that occur frequently in industrial simulations, and quantum algorithms that solve those exact problems either in a shorter time or using less storage than any classical algorithm. We aim for improvements on the scaling with problem-dependent parameters, such as the number of equations to solve, or the desired accuracy. Hence, we consider computations of large mathematical problem size, which are common in simulations.
One common simulation algorithm used in manufacturing is the FEM. This method allows to solve physical equations for continuous systems in return for extensive computational effort. As part of FEM, the physical domain is discretized in so-called elements. The interaction between those elements is described by a system of numerical equations, which often require iterative solvers. For linear equations, both, the computation time ($\mathcal{O}(N^{3})$) and the required storage ($\mathcal{O}(N^{2})$ ) scales polynomially with the number of equations $N$ krishnamoorthy_matrix_2013 . For more details on the finite element method, please refer to Refs. reddy2005 ; huebner1975 .
In order to achieve an advantage, we concentrate on problems that can be solved with exponential speedup over their classical counterparts. Two quantum algorithms promising such a speedup are the HHL algorithm and the QPE, which solve linear equations and eigenvalue problems, respectively. Both problem types can be found in finite element simulations used in manufacturing. Linear equation solvers are more versatile, which makes an end-to-end implementation of the HHL algorithm more favorable. However, since QPE is a significant subroutine of the HHL algorithm, every knowledge gained in an end-to-end implementation of the QPE can be used for the HHL algorithm as well. This is why we concentrate our work on the QPE and eigenvalue problems in QUASIM.
When using QPE, one has to be cautious not to lose any speedup in the subroutines such as the Hamiltonian simulation, state preparation, or post-processing steps. Those subroutines are the interface between classical and quantum computations and scale with the amount of information that needs to enter and exit the quantum computer. Hence, the amount of data that shall be transferred between those two platforms has to be minimal in order to contain quantum speedup ($\lesssim\mathcal{O}(N)$ for polynomial or $\lesssim\mathcal{O}(\log N)$ for exponential speed up). Furthermore, the QPE is a probabilistic algorithm, which means that some results can only be extracted by sampling multiple times from the algorithm.
@@ 3.3 Specific application: milling dynamics simulation
Within the digital twin framework of the Fraunhofer Institute for Productions Technology Aachen dPart® ganser2021 , the engagement simulation is a dexel-based simulation. A dexel model is a discrete representation of a solid volume based on a regular grid, which enables a more efficient simulation of material removal processes than mesh-based methods. For a good representation of 3D volumes, a multi-dexel approximation is used stautner2006simulation . The engagement simulation provides the tool’s contact area, contact length, and width (see Fig. 2). It also allows tracking removed volume to enable an evaluation of the workpiece cut by cut. To compute the removed material, the algorithm calculates the intersections between the milling tool volume and the three-dimensional dexel field. The intersections are performed individually for each of the three dexel grids to take advantage of memory closeness. The calculations performed by the intersectors are based on floating-point calculations.
@@
Starting after the dexel-based engagement simulation in CAM, the main steps of the FEM simulation workflow are: IPW conversion from dexel to solid, meshing process, and modal analysis. Considering the condition for quantum acceleration, the workflow is analyzed step by step both in terms of computational time and with regard to extracting numerical problems that can be transformed in quantum algorithms Schröder24. The “modal analysis”, i.e., solving the eigenvalue problem, is found to be a suitable numerical problem for computation on a quantum computer. The solution of the eigenvalue problem using quantum algorithms is under investigation within this project. Within the classical step “modal analysis” the equation of motion
@@ 4.1 State of the art
Since NISQ computers are still in their infancy, and limited in size and available gate depths, they often can be simulated on classical hardware. Therefore one can use real and simulated NISQ computers alongside classical computers to perform tests on machine learning tasks in hybrid QC. A popular implementation of this uses variational or parameterised quantum circuits (PQCs), which have an advantage over classical ML in certain learning tasks bowles2023contextuality .
@@ 4.2 General approach in QUASIM
Figure 3: The PQC $M_{d}$ used to make temperature predictions in the laser cutting use case. The first qubit, represents $v$ and the subsequent qubits represent the $d$ neighboring nodes. The temperature of each node is embedded on to the corresponding qubit. This is followed by a parameterised layer, where the first qubit and the subsequent qubits are entangled. Finally, the expectation value of the first qubit is measured and then projected to a temperature value.
@@ 4.3 Specific application: laser cutting
Accurately simulating the temperature during laser cutting with complex geometries and long cutting times requires significant computational efforts. In contrast to conventional FEM-based simulations, ML approaches, such as MeshGraphNets pfaff2020mgn , accelerate FEM simulations by one to two orders of magnitude whilst maintaining a high accuracy. However, training of such large ML models currently takes several weeks, which makes updating such models computationally challenging. In this work, the advantages of quantum machine learning approaches, as described above, are investigated to simulate thermodynamics from laser-induced heat on metal plates.
The metal geometry is represented by a mesh structure which can be formulated as a graph $G=(V,E)$. The model is formulated employing the PQC approach outlined in section 4.2 for predicting the temperature $f_{v}^{t+1}$ at the vertex $v$ for the next time step $(t+1)$, depending on the temperatures from the previous time step $t$ at the vertex $v$ and its $d$ neighbors $\mathcal{N}(v)$. A visualisation of the circuit can be seen in Fig. 3.